# Evaluating Generative Models via One-Dimensional Code Distributions

### 09 Mar 2026 v1

@@ Abstract
Most evaluations of generative models rely on feature-distribution metrics such as FID, which operate on continuous recognition features that are explicitly trained to be invariant to appearance variations, and thus discard cues critical for perceptual quality. We instead evaluate models in the space of discrete visual tokens, where modern 1D image tokenizers compactly encode both semantic and perceptual information and quality manifests as predictable token statistics. We introduce Codebook Histogram Distance (CHD), a training-free distribution metric in token space, and Code Mixture Model Score (CMMS), a no-reference quality metric learned from synthetic degradations of token sequences. To stress-test metrics under broad distribution shifts, we further propose VisForm, a benchmark of 210K images spanning 62 visual forms and 12 generative models with expert annotations. Across AGIQA, HPDv2/3, and VisForm, our token-based metrics achieve state-of-the-art correlation with human judgments, and we will release all code and datasets to facilitate future research.
@@ 6 Conclusion
We introduce a discrete token-based paradigm for evaluating generative models by shifting from continuous features to codebook statistics. This framework enables two complementary metrics: CHD for distribution matching and CMMS for reference-free quality assessment. Both achieve state-of-the-art correlation with human judgment across multiple benchmarks, including our VisForm dataset. Our approach is scalable, interpretable, and robust to domain shifts, establishing a unified framework for perceptually aligned quality assessment. Future work includes modeling higher-order token statistics for better spatial structure capture and extending to video and 3D generation evaluation.

### 12 Mar 2026 v2

@@ Abstract
Most evaluations of generative models rely on feature-distribution metrics such as FID, which operate on continuous recognition features that are explicitly trained to be invariant to appearance variations, and thus discard cues critical for perceptual quality. We instead evaluate models in the space of discrete visual tokens, where modern 1D image tokenizers compactly encode both semantic and perceptual information and quality manifests as predictable token statistics. We introduce Codebook Histogram Distance (CHD), a training-free distribution metric in token space, and Code Mixture Model Score (CMMS), a no-reference quality metric learned from synthetic degradations of token sequences. To stress-test metrics under broad distribution shifts, we further propose VisForm, a benchmark of 210K images spanning 62 visual forms and 12 generative models with expert annotations. Across AGIQA, HPDv2/3, and VisForm, our token-based metrics achieve state-of-the-art correlation with human judgments. We will release all code and datasets to facilitate future research, with the code publicly available at https://github.com/zexiJia/1d-Distance.
@@ 6 Conclusion
We present a discrete token-based framework for generative model evaluation by shifting from continuous features to codebook statistics. Based on this paradigm, we develop two complementary metrics: CHD for distribution matching and CMMS for reference-free quality assessment. Both achieve state-of-the-art correlation with human judgment across multiple benchmarks, including VisForm. Our framework is scalable, interpretable, and robust to domain shifts, offering a unified solution for perceptually aligned quality assessment. Future work will explore higher-order token statistics for improved spatial modeling, as well as extensions to video and 3D generation.
Acknowledgements
This work was completed at the Pattern Recognition Center, WeChat AI, Tencent. The authors would like to express their sincere gratitude to Professor Anil K. Jain for his valuable guidance, insightful discussions, and continuous encouragement, which greatly benefited this work.